OpenStudy (anonymous):

PLEASE ITS MY LAST QUESTION IM TIRED Solve four x is less than or equal to negative twenty for x. Show your work.

OpenStudy (anonymous):

Just write down the equation, 4x>(oet)-20 Treat the GTOET as a equal sign instead of GTOET solve, and you get -5.

OpenStudy (anonymous):

\[4x \ge -20\]

OpenStudy (anonymous):

put a line through the ge sign and subtract 20 on each side.

OpenStudy (anonymous):

so what is x?

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

cause -5 times 4 equals 20

OpenStudy (texaschic101):

nicely done :)

OpenStudy (whpalmer4):

\[4x \le -20\]Divide both sides by 4\[\frac{4x}{4} \le \frac{-20}{4}\]\[x \le -5\] Thing to be careful about is that if you multiply or divide by a negative number, you need to flip the inequality sign to point the other way!